OpenStudy (anonymous):

How will I know if the data that I gathered is accurate using Type of Waveform [RMS-Value]/[peak-value] Sinusoid [2]^-0.5 Square 1 Triangular [3]^-0.5

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

Not clear exactly what you are asking. You can identify the peak value easily. You can also square the values and find the mean of the squares and then take the square root of that to get the root mean square value. The shapes of these waveforms are so different that you are not likely to need this metric to determine which shape you have. Good night.

OpenStudy (anonymous):

I have gathered the frequency and voltage for each shape and I need to use the rms/peak to determine whether my collected data is accurate...

OpenStudy (anonymous):

ok.. there is a particular relationship between peak value and rms value.. for every type of wave.. its always a constant for a given wave.. so you need to evaluate it theoretically and check your answer!

OpenStudy (anonymous):

@Mashy How will I do that?

OpenStudy (anonymous):

I have frequency too

OpenStudy (anonymous):

By using the definition and integration.. (hope you have taken calculus) Peak value is simple.. any Ac waveform is given in terms of its peak value.. Rms value .. = root of mean of square of the values so \[RMS = \sqrt(\frac{\int\limits_{0}^{T}(f(x))^2}{T})\]

OpenStudy (anonymous):

I took calculus, The only problem is that I am new to circuit analysis, as in new. I have the peak( i got it from an oscilloscope)... the f(x)...what is it? I am assuming that T is time. But I did not get the time

OpenStudy (anonymous):

i am not sure whether this is what is expected out of you but every wave can be written in terms of a equation sinwave for example f(x) = V0sin(wt) so V0 is the peak value, and w = 2*pi/f gives you angular frequency.. and then u can find RMS using the above integral mentioned..

OpenStudy (anonymous):

T = time period = 1/f

OpenStudy (anonymous):

it should be f(t) sorry :P

OpenStudy (anonymous):

How about if it is not a sinusoidal function i.e. a triangular, or square will I still use the same formula?

OpenStudy (anonymous):

Yes.. that is the most general formula applicable for any periodic function!

OpenStudy (anonymous):

thank you